OpenStudy (rayep12):

Philip is going on a 4000-kilometer road trip with three friends. The car consumes 6 liters of gas per 100 kilometers, and gas costs $1.50 per liter

OpenStudy (rayep12):

If Philip and his friends want to split the cost of gas evenly, how much should they each pay?

OpenStudy (rayep12):

sorry reset question

OpenStudy (yanasidlinskiy):

\[(4000)* \frac{ 6liters }{ 100km } * \frac{ $1.50 }{ liters } = ?\]

OpenStudy (whpalmer4):

they ought to pay for a tuneup, instead! 66 liters of fuel to go 100 km?!? Are you sure you copied the question correctly?

OpenStudy (agent0smith):

Didn't you just do one exactly like this? You shouldn't need to have someone walk you through it again.

OpenStudy (yanasidlinskiy):

LOL. You better have hahha

OpenStudy (whpalmer4):

\[4000 \text{ km}*\frac{6.6 \text{ liter}}{100 \text{ km}}*\frac{\$1.50}{1\text{ liter}}=\] (I am assuming you dropped a decimal point from the liters per km figure) That will give you the total fuel cost. How much will each person pay?